Use the quadratic ,x²+18x=25 to answer the following questions.part A:find the value of c that will result in a perfect square trinomial.x²+18x+c=25+c. Part B:rewrite the above equations into a perfect square binomial form . (x+—)²=— part c:what is the value of x?
Answer by Boreal(15235)

x²+18x=25
x^2+18x+81=25+81, because half of the 18 squared will give a perfect square. Add 81 to both sides
(x+9)^2=106 There is now a perfect square on the left, take the square root of it and move the constant over to the right. Can check using the quadratic formula.
x+9=sqrt (106)
x=-9+/- sqrt (106)
-------------------------